Countdown: Days-2 1ime:-10:17:03 Question 11 The distribution of IQ (Intelligence Quotient) is approximately normal inType numbers in the boxes. shape with a mean of 100 and a standard deviation of 12. o points According to the standard deviation rule, % of people have an IQ between 76 and 124. Do not round. Question 12 Type numbers in the boxes. The distribution of IQ (Intelligence Quotient) is approximately normal in shape with a mean of 100 and a standard deviation of 19 1o points According to the standard deviation rule, only % of people have an IQ over 119. Question 13 numbers in the boses The distribution of the amount of money spent by students on textbooks in a ype semester is approximately normal in shape with a mean of: u 421 and a 1o points standard deviation of:o# 22. According tothe standard deviation rule, almost 2.5% of the students spent more than what amount of money on textbooks in a semester?
In June 2017, a survey was conducted in which a random sample of 1444 U.S. Select one answe adults was asked the following question: zo points In 1973 the Roe versus Wade decision established a womans constitutional right to an abo least in the first three months of pregnancy. Would you like to see the Supreme Court compl overturn its Roe versus Wade decision, or not? The results were: Yes-2696, No-67%, Unsure-7% Which of the following is true about this scenario? A. 26%, 67%, and 7% are all parameters. B. 26%, 67%, and 796 are all statistics. C. O If another random sample of size 1444 U.S. adults were to be chosen, we woul get the exact same distribution of answers. Question 10 Type numbers i The distribution of the amount of money spent by students on textbooks in a semester is approximately normal in shape with a mean of 324 and a standard |ツPart 13 pa deviation of 25 According to,the standard deviation rule, approximately 99.7% of the students spent betweens ands on textbooks in a semester. by Part 2:
Answer 10-13 first please
Answer #1

The given problem is solved using the concept of empirical rule or standard deviation rule.

Then 서 to Heee 324 t 3x25, 37975 ση tert books un, a keulaa. 2 (X >411+2x22) = P(X> 465)= o-25 Ό. o-025-2.5% or 2, 57.